OpenStudy (anonymous):

What is the Total Energy of an object that has mass of 6.98 kg, with a velocity of 1.92 m/s, which is also experiencing an acceleration of 9.94 m/s2 at a height of 9.43 m?

OpenStudy (anonymous):

((1/2)mv^2)+mah=E We're given m, v, a, AND h--that's literally every variable with which we can solve for E.

OpenStudy (anonymous):

That's what I thought, the problem is that the problem icludes an acceleration component: 9.94 m/s^2. That means that: E=KE + PE KE = 1/2m * v^2 PE = mgh So it would seem pretty simple: KE = 1/2 * m * v^2 KE = 1/2 * 6.98kg * (1.92) ^2 KE = 12.86 J (Joules) PE = mgh PE = 6.98 * 9.81 * 9.43 PE = 645.7 J Therefore, E = PE + KE E = 645.7 J + 12.86 J E = 658.57 J But that answer is WRONG..... it's that extra A = 9.94 m/s^2 that is throwing me off somewhere...

OpenStudy (anonymous):

I also replaced the 9.81 (g) value with the 9.94 (A) value but the new answer (667.12 J) is still wrong.

OpenStudy (anonymous):

Interesting. Energy is a scalar, so I don't see there being any vector issues. Give me some time to brood over my mistake. XD

OpenStudy (anonymous):

thanks!